OpenStudy (anonymous):

How many solutions does the following equation have? ln(x^{2} + 4x − 5) = 0

zepdrix (zepdrix):

\[\large \ln\left(x^2+4x-5\right)=0\] Hmm ok this problem is a little strange. Here is one way to approach it at least. Let's exponentiate both sides, (rewrite both sides with a base of e) \[\huge e^{\ln\left(x^2+4x-5\right)}=e^0\] Since the exponential and the logarithm are inverse operations of one another, they essentially cancel out. Giving us,\[\large x^2+4x-5=e^0\]e^0 is just 1.\[\large x^2+4x-5=1\]Subtract 1 from each side,\[\large x^2+4x-6=0\] If you throw this into the Quadratic Formula you'll be able to find out how many solutions exist. :D If that's way too confusing you can let me know. Maybe there is another way you're suppose to approach this.

OpenStudy (anonymous):

oh my gosh @zepdrix i get it =P that is strange